OpenStudy (akgorgeous):

CHECK MY WORK- SYSTEM OF EQUATIONS- 2x + y = 5 x − 2y = 10 y = 5 − 2x x − 2(5 − 2x) = 10 x + 10 + 4x = 10 5x + 10 = 10 5x = 0 x = 0 2(0) + y = 5 y = 5

OpenStudy (akgorgeous):

@mathmale help?

OpenStudy (mathmale):

You have found (0,5) as a possible solution. Subst. these x- and y-coordinates into both of the given equations, one equation at a time, to determine whether or not the equation is then true.

OpenStudy (alivejeremy):

Mathmale 💎 - 💎

OpenStudy (mathmale):

If both equations are true, what does that tell you?

OpenStudy (akgorgeous):

there is no error?

OpenStudy (akgorgeous):

@steve816

OpenStudy (triciaal):

substitution ok error on line 3 -2 * 5 = -10

OpenStudy (mrnood):

if x=0 and y=5 what is x-2y ? Is this compatible with the question?

OpenStudy (triciaal):

with x = 4 y = 5 - 2*4 = -3 check in equation 2 left x - 2y = 4-2*-3 = 4 + 6 = 10 =right so correct (4, -3)

OpenStudy (akgorgeous):

So I combined like terms incorrectly, it should have been 4x instead of 5x

OpenStudy (triciaal):

no

OpenStudy (triciaal):